Find the cube root of each of the following numbers by prime factorisation method.

 

(iv) 27000

 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


(iv) 27000

Solution:

27000 = 2×2×2×3×3×3×3×5×5×5

By grouping the factors in triplets of equal factors, 27000 = (2×2×2)×(3×3×3)×(5×5×5)

Here, 27000 can be grouped into triplets of equal factors,

∴ 27000 = (2×3×5) = 30

Hence, 30 is cube root of 27000.



Thanks